# Benchmark the Linguistic Competence of Language Models

### v1

@@ Abstract
We introduce Holmes, a benchmark to assess the linguistic competence of language models (LMs) – their ability to grasp linguistic phenomena. Unlike prior prompting-based evaluations, Holmes assesses the linguistic competence of LMs via their internal representations using classifier-based probing. In doing so, we disentangle specific phenomena (e.g., part-of-speech of words) from other cognitive abilities, like following textual instructions, and meet recent calls to assess LMs’ linguistic competence in isolation. Composing Holmes, we review over 250 probing studies and feature more than 200 datasets to assess syntax, morphology, semantics, reasoning, and discourse phenomena. Analyzing over 50 LMs reveals that, aligned with known trends, their linguistic competence correlates with model size. However, surprisingly, model architecture and instruction tuning also significantly influence performance, particularly in morphology and syntax. Finally, we propose FlashHolmes, a streamlined version of Holmes designed to lower the high computation load while maintaining high-ranking precision.
@@ 1 Introduction
Linguistic competence is the unconscious understanding of language, like grasping grammatical rules (Chomsky, 1965). As language models (LMs) are trained on simple tasks like next word prediction (Brown et al., 2020), one might naturally wonder: What is the linguistic competence of LMs, and how do they differ? To answer such questions, benchmarks estimate cognitive abilities by providing textual instructions and evaluate LMs’ responses, as done for mathematical reasoning (Cobbe et al., 2021) or factual knowledge (Petroni et al., 2019, 2020). However, they conflate latent abilities (like following provided instructions) with those under test, such as understanding specific linguistic phenomena, e.g., syntactic structures (Liang et al., 2023). As this entanglement makes it infeasible to draw definitive conclusions about distinct abilities (Hu and Levy, 2023), recent studies call to assess the linguistic competence of LMs comprehensively and in isolation (Lu et al., 2023; Mahowald et al., 2024).
@@
In this work, we introduce the Holmes (Figure 2). A benchmark to assess the linguistic competence of LMs (Figure 1) regarding numerous linguistic phenomena. To fully disentangle the understanding of these phenomena and other abilities of LMs, we use classifier-based probing (Tenney et al., 2019a; Hewitt and Manning, 2019; Belinkov, 2022). A method that uses the LMs’ internal representations of text inputs to train linear models (probes) to predict specific aspects of phenomena, such as words’ part-of-speech (POS). We then approximate the LMs’ grasp of these phenomena using the probes’ performance, rigorously verified using control tasks (Hewitt and Liang, 2019) and from an information theory perspective (Voita and Titov, 2020). With this particular and comprehensive scope, we thoroughly address the initially raised questions as follows:
@@
The review of over 270 probing studies reveals a gap in comprehensively evaluating linguistic competence. Despite covering over 200 probing tasks and 150 LMs, individual studies focus on particular tasks and LMs. As a result, only three LMs were probed on over 20% of the tasks, and one single task was evaluated for more than 20% of the reviewed LMs. Notably, recent large LMs are significantly underrepresented.
@@
To address this identified deficiency, Holmes offers a structured framework to assess the English linguistic competence of LMs comprehensively. It features 208 distinct datasets covering morphology, syntax, semantics, reasoning, and discourse phenomena, including previously underrepresented ones like negation or rhetoric in text (Liang et al., 2023).
@@
From assessing 59 LMs (Figure 1), we find that no single one consistently excels the others and that their linguistic competence is more pronounced for morphology and syntax than the other phenomena types. Instead, we find model size, model architecture, and instruction tuning fundamentally affect their linguistic competence.
First, LMs’ linguistic competence, particularly for morphology and syntax, scales with their model size. This generalizes previous findings (Tenney et al., 2019b; Zhang et al., 2021) beyond LMs with 350 million parameters. Second, contrary to prompting evaluations (Lu et al., 2023) and aligned with other work (Waldis et al., 2024a; Gautam et al., 2024), model architecture is critical. The linguistic competence of decoder-only LMs is less pronounced, and even 70 billion does not allow them to encode linguistic phenomena of words with comparable strength to encoder-only LMs of a similar size. Third, while previous studies focused on aligning LMs with human interactions through instruction tuning (Ouyang et al., 2022; Touvron et al., 2023; Zhou et al., 2023), we show for the first time its effect on their linguistic competence. It improves morphology and syntax but has mixed effects for the other types of phenomena. Lastly, we contrast the results of Holmes with OpenLLM (Beeching et al., 2023), an extensive LM benchmark focusing on user-centered applications like mathematical reasoning. We find that Holmes provides a unique but supplementary perspective, as rankings partly align, especially for reasoning-related phenomena.
@@
We summarize our contributions as follows:
  • Benchmark. Holmes comprehensively and thoroughly assesses the linguistic competence of LMs in isolation, providing substantial ground for advancements in NLP.
  • Empirical insights. Extensive experiments reveal that LMs’ linguistic competence is more pronounced for morphology and syntax, and size, architecture, and instruction tuning are crucial for LM differences.
  • Ease of use. We provide tools to interactively explore Holmes results and straightforward code to evaluate upcoming LMs with efficiency in mind (FlashHolmes).
@@ 2 Preliminaries
Language Models compute probabilities for word sequences $i$, enabling tasks such as classifying $i$, textual comparisons between $i$ and another sequence $i^{\prime}$, and text generation based on $i$. We consider LMs as any model producing representations of input $i$, regardless of their specific type: sparse like bag-of-words (Harris, 1954); static such as GloVe (Pennington et al., 2014); or contextualized transformer-based LMs (Devlin et al., 2019; Raffel et al., 2020).
Linguistic Competence
Following Chomsky (1965), linguistic competence is defined as the unconscious knowledge of language, encompassing the understanding of specific linguistic phenomena, including word dependencies and their distinct parts of speech (POS).
@@
We define a dataset as text examples and labels covering a specific aspect of a linguistic phenomenon, like words and their POS tag. Typically, these labels are highly unambiguous to assess the specific aspect under test in isolation.
@@
Using probes, we empirically assess the linguistic competence of LMs regarding the featured linguistic phenomena in Holmes. To this end, we employ probing tasks using the widely recognized classifier-based probing method (Tenney et al., 2019a; Hewitt and Manning, 2019; Belinkov, 2022), or known as diagnostic classifiers (Veldhoen et al., 2016; Giulianelli et al., 2018). Running such a probing task involves training a probe (linear model) using the specific dataset to test a distinct aspect of a linguistic phenomenon in isolation. Therefore, we feed the text examples, encoded with a given LM, as training inputs. Subsequently, we use the probe’s performance to approximate how an LM understands the specific linguistic phenomenon under test. With a higher score, we assume the embeddings embody patterns relevant to this phenomenon, which enhances the accuracy (Tenney et al., 2019b).
@@ 3 Meta-Study
In this section, we survey 274 studies (§ 3.1), probing LMs’ linguistic competence. We analyze these studies regarding their evolution, covered probing tasks and LMs (§ 3.2), and identify the apparent need for consolidating existing resources (§ 3.3).
@@ 3.1 Scope
We analyze 28k papers ($P$) from 2015 to August 2023 of major NLP conferences (TACL, ACL, AACL, COLING, EACL, EMNLP, NAACL, and corresponding workshops) expanded with selected work from other venues such as ICLR. To identify relevant work, we follow a semiautomatic approach. First, we automatically select papers based on their meta-data and full text. [^1] We use PyPDF2 v3.0.0, DBLP and semanticscholar API. We select a total of 493 candidate papers matching at least one of the following three criteria ($P^{\prime}=\{\forall p\in P|p\in P_{1}\cup p\in P_{2}\ \cup p\in P_{3}\}$):
$\mathbf{P_{1}}$: papers contain probing or probe in the title.
$\mathbf{P_{2}}$: papers contain probing or probe in the abstract and at least five times in the main content.
$\mathbf{P_{3}}$: papers contain probing or probe at least ten times in the main content.
We manually verified these automatically curated candidates ($P^{\prime}$) and found 274 relevant papers ($P_{r}$).We selected them as they either evaluate LMs regarding one or more linguistic phenomena as part of the analysis or as a main contribution. This involves filtering papers using the term probing in other senses, such as probing hash tables (Bogoychev and Lopez, 2016).
@@ 3.2 Analysis
First, we analyze the evolution of the relevant studies. Figure 3 relates how these studies cite each other (probing citations $C_{p}$) compared to other gathered citations (general citations $C_{g}$). Colorized, we show the ratio $\alpha$ between these two measures $\alpha=\frac{|C_{p}|+1}{|C_{g}|+1}$. First, only a fraction of the works gained general attention, as 16 papers exceeded 200 general citations. Further, probing works cite each other rather sparsely, with an average probing citation ratio of $\alpha=0.1$. Therefore, we see other fields are paying little attention to the linguistic competence of LMs. Paired with scattered citation patterns, we identify the need to consolidate existing resources to solidly ground research in this field.
@@
We categorize the selected work according to their probing focus: methodological, new methods, like control tasks (Hewitt and Liang, 2019) or minimum description length (Voita and Titov, 2020); task-focused assessing specific linguistic phenomena as main contributions, such as discourse relations in text (Koto et al., 2021); and analytical using probing tasks to analyze LMs, such as the impact of pre-training data (Zhang et al., 2021). Figure 4 shows: the majority (51.8%) of studies focus on specific probing tasks like numeric scales (Zhang et al., 2020), or morphosyntactic (Shapiro et al., 2021); 35.7% use probing as a supplementary analytical tool, for example, analyzing the effect of fine-tuning (Mosbach et al., 2020a; Zhu et al., 2022a); 12.5% address methodological problems related to probing (Wu et al., 2020; Immer et al., 2022; Zhu et al., 2022b).
@@
Next, we analyze the specific employed probing method: classifier, using linear or shallow models to probe internal representations of LMs, as demonstrated in Tenney et al. (2019a); mask, letting LMs fill gaps to verify linguistic phenomena, as shown in Talmor et al. (2020) or Warstadt et al. (2020); attention, which relies on attention patterns, as used in Pandit and Hou (2021) for bridging; and other, methods not belonging to the previous three categories, such as dimension selection (Torroba Hennigen et al., 2020). Most studies utilize the classifier-based probing method (74%), 20% conduct mask-based probing, and only a minority of work ($\sim$ 3%) considers attention patterns or other approaches.
@@
Finally, we analyze which tasks and LMs the relevant probing studies consider. For example, Tenney et al. (2019b) considers BERT and probes POS tagging, semantic-role labeling (SRL), and other ones. Aggregated over all studies, we found a broad coverage of 289 unique tasks and 161 distinct LMs. Below, we delve into the details and highlight noteworthy findings.
We analyze how LMs and tasks are considered jointly in Figure 5. Despite the broad coverage, single studies, including fundamental ones, maintain a particular focus and consider only a fraction of LMs and tasks. For example, while most tasks (72%) were assessed on BERT, RoBERTa’s coverage has already declined to 42%. Conversely, part-of-speech tagging (POS), the most probed task, was only evaluated on 23% of the LMs, for example, not covering prominent examples like BART Lewis et al. (2020). Notably, more recently released larger and powerful LMs, like PYTHIA (Biderman et al., 2023), UL2 (Tay et al., 2023), or LLAMA-2 (Touvron et al., 2023), and instruction-tuned LMs (FLAN-T5 (Chung et al., 2022), LLAMA-2-Chat (Touvron et al., 2023), or TK-Instruct (Wang et al., 2022) are missing almost entirely, with single more recent exceptions (Hu and Levy, 2023; Waldis et al., 2024a). Again, these insights underscore the need to consolidate existing resources for more dense coverage. This is further evident when considering Figure 5, where we sort LMs and tasks according to how often they were mentioned in the relevant works. Then, we plot their cumulative coverage concerning all mentions. For example, considering all studies (red line), the top-10 most mentioned LMs account for 80% of all LMs mentions (black dot). In contrast, the other 151 unique LMs account for only 40%. Comparing the paper focus, we see that methodological studies rely only on 24 LMs and 36 tasks. In contrast, task-focused and analytical work covers a similar number of LMs (91 and 99, respectively). However, due to their distinct focus, task-focused studies cover significantly more tasks (202) than analytical ones (115).

### v4

# A Benchmark to Assess the Linguistic Competence of Language Models

@@ Abstract
We introduce Holmes, a new benchmark designed to assess language models (LMs) linguistic competence - their unconscious understanding of linguistic phenomena. Specifically, we use classifier-based probing to examine LMs’ internal representations regarding distinct linguistic phenomena (e.g., part-of-speech tagging). As a result, we meet recent calls to disentangle LMs’ linguistic competence from other cognitive abilities, such as following instructions in prompting-based evaluations. Composing Holmes, we review over 270 probing studies and include more than 200 datasets to assess syntax, morphology, semantics, reasoning, and discourse phenomena. Analyzing over 50 LMs reveals that, aligned with known trends, their linguistic competence correlates with model size. However, surprisingly, model architecture and instruction tuning also significantly influence performance, particularly in morphology and syntax. Finally, we propose FlashHolmes, a streamlined version that reduces the computation load while maintaining high-ranking precision.
@@ 1 Introduction
Linguistic competence is the unconscious understanding of language (Chomsky, 1965), like the syntactic structure of a sentence. As language models (LMs) are trained on simple tasks such as next word prediction (Brown et al., 2020), one might naturally wonder: What is the linguistic competence of LMs, and how do they differ? To answer such questions, contemporary benchmarks estimate cognitive abilities, as done for mathematical reasoning (Cobbe et al., 2021) or factual knowledge (Petroni et al., 2019b, 2020). However, such benchmarks rely on LMs’ use of language (textual responses) known as linguistic performance (Matthews, 2014). As a result, they conflate abilities tested with specific instructions, as done for syntactic phenomena in Blevins et al. (2023a), with latent abilities like producing coherent text or following instructions. As this entanglement makes it infeasible to draw definitive conclusions (Hu and Levy, 2023; Liang et al., 2023; Perlitz et al., 2024), recent studies call to assess LMs’ linguistic competence comprehensively and isolated (Lu et al., 2023; Mahowald et al., 2024).
Figure 1: In Holmes, we encode examples of probing datasets using frozen LMs. Then, we train probes (linear models) with labels representing the specific linguistic phenomenon under test. Finally, we use the results of testing the probes to approximate the LMs’ linguistic competence regarding the tested phenomena.
@@
In this work, we introduce Holmes (Figure 2), a benchmark to assess the linguistic competence of LMs (Figure 7) regarding numerous linguistic phenomena. To disentangle LMs understanding of these phenomena from their linguistic performance, we assess the LMs’ internals using classifier-based probing (Tenney et al., 2019a; Hewitt and Manning, 2019; Belinkov, 2022). As illustrated in Figure 1 for probing the part-of-speech (POS) tags for words, we first train linear models (probes) using the internal representations of text inputs from the last model layer to predict the specific phenomena aspects. We then approximate the LMs’ grasp of these phenomena using the probes’ performance, rigorously verified using control tasks (Hewitt and Liang, 2019) and from an information theory perspective (Voita and Titov, 2020). With this particular and comprehensive scope, we thoroughly address the initially raised questions as follows:
@@
The review of over 270 probing studies reveals a gap in comprehensively evaluating linguistic competence. Despite covering over 200 probing tasks and 150 LMs, individual studies focus on particular tasks and LMs. As a result, only three LMs were probed on over 20% of the tasks, and only one task (POS) was evaluated for more than 20% of the LMs. Notably, recent large LMs are significantly underrepresented.
@@
Addressing these identified deficiencies, Holmes offers a structured way to assess LMs’ English linguistic competence comprehensively. It features 208 distinct datasets covering morphology, syntax, semantics, reasoning, and discourse phenomena, including previously underrepresented ones like negation or rhetoric.
@@
From assessing 59 LMs, we find that no LM consistently excels the others. Further, linguistic competence is more pronounced for morphology and syntax than the other types of phenomena, and LMs’ linguistic competence is fundamentally affected by model size, model architecture, and instruction tuning.
First, we generalize previous findings (Tenney et al., 2019b; Zhang et al., 2021) and show LMs’ linguistic competence, particularly morphology and syntax, scales beyond 350 million parameters. Second, contrary to the prompting evaluations (Lu et al., 2023) and aligned with Waldis et al. (2024a) and Gautam et al. (2024), model architecture is critical. The linguistic competence of decoder-only LMs lags behind encoder-only ones. Not even 70 billion decoders produce representations for words with the same stability as encoders with 110 million parameters. Third, while instruction tuning (Ouyang et al., 2022; Touvron et al., 2023; Zhou et al., 2023) aims to align LMs with human interactions, we focus for the first time on its effect on linguistic competence. We found that instruction tuning improves morphology and syntax but has mixed effects on other phenomena types, hinting at a superficial alignment. Lastly, we compare Holmes with other benchmarks. While LM rankings of reasoning-intense downstream tasks (Beeching et al., 2023) correlate with reasoning phenomena, explicitly prompting for linguistic phenomena (Liang et al., 2023) leads to unreliable results. As these results show that Holmes aligns with other benchmarks, its probing-based evaluation is indispensable for explicitly testing LMs’ linguistic competence disentangled from their linguistic performance.
@@
Contributions
With Holmes, we introduce a comprehensive and thorough benchmark to assess LMs’ linguistic competence, providing ground to evaluate them more holistically. Extensive experiments on Holmes reveal that LMs’ linguistic competence is manifold and more pronounced for phenomena targeting words and syntactic structure than semantic, reasoning, or discourse. LMs properties like size or architecture crucially account for differences among LMs. Fostering further research, we provide interactive tools to explore Holmes and straightforward evaluation code for upcoming LMs with efficiency in mind.
@@ 2 Preliminaries
Language Models compute probabilities for word sequences $i$, enabling tasks such as classifying $i$, textual comparisons between $i$ and another sequence $i^{\prime}$, and text generation based on $i$. We consider LMs as any model producing representations of $i$, regardless of their specific type: sparse like bag-of-words (Harris, 1954); static such as GloVe (Pennington et al., 2014); or contextualized transformers (Devlin et al., 2019; Raffel et al., 2020).
Linguistic Competence and Performance
For centuries (Robins, 2013), linguists have been fascinated by the processes of language learning, usage, and evolution. One specific discussion is the differentiation between knowing and using a language. de Saussure (1916) distinguished between language with specific rules and words (langue) as an ongoing negotiated fulfillment of the societal need for communication and its usage (parole). Similarly, Chomsky (1965) uses the term linguistic competence for the unconscious understanding of language and linguistic performance for using languages in any utterance. In this work, we follow Chomsky’s terminology and treat LMs as static artifacts of a certain time, omitting ongoing processes of the society considered by de Saussure. Specifically, we focus on assessing the linguistic competence of LMs, including specific linguistic phenomena like word dependencies and their distinct parts of speech (POS). Opposed, contemporary benchmarks (Cobbe et al., 2021; Petroni et al., 2019b, 2020) assess linguistic performance by providing textual instructions and verifying LMs’ textual responses. Note that this evaluation protocol can also verify an understanding of specific linguistic phenomena, as done in Blevins et al. (2023a) or Liang et al. (2023) for syntactic structure. However, such evaluation protocols conflate LMs’ linguistic competence with latent abilities (like following instructions. Thus, Holmes unique evaluation perspective is indispensable to assess linguistic phenomena isolated to assess LMs comprehensively.
@@
We define a dataset as text examples and labels covering a specific aspect of a linguistic phenomenon, like words and their POS tags. Typically, these labels are unambiguous, enabling us to assess the specific aspect under test in isolation.
@@
Using probes, we empirically assess the linguistic competence of LMs regarding the featured linguistic phenomena in Holmes. We design probing tasks using the widely recognized classifier-based probing method (Tenney et al., 2019a; Hewitt and Manning, 2019; Belinkov, 2022) also known as diagnostic classifiers (Veldhoen et al., 2016; Giulianelli et al., 2018). Running such a probing task involves training a probe (linear model) using the specific dataset to test a distinct aspect of a linguistic phenomenon in isolation. To do this, we encode the text examples of a dataset with a given LM and use them to train the probe regarding the specific labels representing the tested linguistic phenomenon. The probe’s performance is then used to approximate the LM’s understanding of the specific phenomenon. A higher score indicates that LMs capture patterns relevant to this phenomenon internally, which in turn enhances the accuracy (Tenney et al., 2019b).
@@ 3 Meta-Study
This section summarizes our survey of 274 studies (§ 3.1) probing LMs’ linguistic competence. We analyze them regarding their evolution, probing tasks and LMs addressed (§ 3.2), and identify the need to consolidate existing resources (§ 3.3).
@@ 3.1 Scope
We analyze 28k papers ($P$) from 2015 to August 2023 of major NLP conferences (TACL, ACL, AACL, COLING, EACL, EMNLP, NAACL, and the corresponding workshops) expanded with selected work from other venues such as ICLR. To identify relevant work, we employ a semi-automatic approach. First, we use automated filtering based on paper metadata and full text [^1] We use PyPDF2 v3.0.0, DBLP and semanticscholar API., grounded in the occurrence of established terminology related to the specific focus of Holmes, namely disentangling the linguistic competence of LMs by studying their internal representations. This terminology, including probing and probe, is commonly found in influential literature surveys (Rogers et al., 2020; Belinkov, 2022) and diverse investigation settings, such as analyzing internal representations using linear classifiers (Tenney et al., 2019b; Conneau et al., 2018; Elazar et al., 2021) or masked-based approaches focusing on lexical knowledge of LMs (Petroni et al., 2019a; Talmor et al., 2020a; Kassner et al., 2021; Peng et al., 2022). Specifically, we define three criteria to identify relevant papers: $P^{\prime}=\{\forall p\in P|p\in P_{1}\cup p\in P_{2}\ \cup p\in P_{3}\}$, where:
$\mathbf{P_{1}}$: papers with probing or probe in the title.
$\mathbf{P_{2}}$: papers with probing or probe in the abstract and at least five occurrences in the main content.
$\mathbf{P_{3}}$: papers with probing or probe occurring at least ten times in the main content.
We identified 493 matching papers ($P^{\prime}$) by applying these criteria. We then manually review the automatically generated candidate list ($P^{\prime}$) and select studies that examined LMs with one or more specific linguistic phenomena as part of their analysis or as a primary contribution. This process involves filtering out papers using the term probing in other senses, such as probing hash tables in Bogoychev and Lopez (2016). Moreover, we supplement the candidates with a curated selection of highly relevant studies that do not meet the above criteria. For example, seminal works published before 2019 which employ terms like “diagnostic classifier(Giulianelli et al., 2018; Hupkes and Zuidema, 2018), as well as other notable studies (Gupta et al., 2015; Shi et al., 2016). This comprehensive approach yields 274 relevant papers ($P_{r}$), which we further analyze subsequently.
@@ 3.2 Analysis
We begin by examining the evolution of relevant studies in the field, illustrated in Figure 3. We analyze the citation patterns among these studies, distinguishing between probing citations ($C_{p}$), which represent citations between them, and general citations ($C_{g}$), which encompass all other citations. The colorized ratio $\alpha=\frac{|C_{p}|+1}{|C_{g}|+1}$ visually relates these two measures. This analysis reveals that only a small fraction of the works have garnered broad recognition, with 16 papers exceeding 200 general citations. Furthermore, probing works cite each other relatively infrequently, with an average probing citation ratio of $\alpha=0.1$. This suggests that other fields have paid limited attention to LMs’ linguistic competence. The scattered citation patterns and lack of engagement with this topic underscore the need to consolidate existing resources and establish a solid foundation to bootstrap research in this area.
@@
We categorize the selected work according to their probing focus into three categories: methodological, which introduces new methods, such as control tasks (Hewitt and Liang, 2019) or minimum description length (Voita and Titov, 2020); task-focused, which assesses specific linguistic phenomena as main contributions, such as discourse relations in text (Koto et al., 2021); and analytical, which uses probing tasks to analyze LMs, such as the impact of pre-training data (Zhang et al., 2021). As shown in Figure 4, the majority of studies (51.8%) focus on specific probing tasks, such as numeric scales (Zhang et al., 2020), or morphosyntactic analysis (Shapiro et al., 2021). A significant proportion (35.7%) use probing as a supplementary analytical tool, for example, to analyze the effect of fine-tuning (Mosbach et al., 2020a; Zhu et al., 2022a). The remaining 12.5% address methodological problems related to probing (Wu et al., 2020; Immer et al., 2022; Zhu et al., 2022b).
@@
Next, we analyze the specific employed probing method regarding four categories: (1) classifier-based probing, which uses linear or shallow models to probe internal representations of LMs; (2) mask-based probing, where LMs fill gaps to verify linguistic phenomena; (3) attention-based probing, which relies on attention patterns; and (4) other methods that do not fit into the previous three categories. Our analysis indicates that most studies (74%) utilize the classifier-based probing method, as exemplified in Tenney et al. (2019a). Additionally, 20% of studies conduct mask-based probing, as shown in Talmor et al. (2020b). In contrast, only a small portion of work ($\sim$ 3%) considers attention patterns or other approaches, such as bridging (Pandit and Hou, 2021) or dimension selection (Torroba Hennigen et al., 2020).
@@
Finally, we examine the tasks and LMs investigated by the relevant studies. For example, Tenney et al. (2019b) explore BERT on various tasks, including POS tagging, semantic-role labeling (SRL), and others. Our analysis reveals that, collectively, these studies cover a remarkable 289 unique tasks and 161 distinct LMs, demonstrating a broad scope of investigation. Below, we delve into the details and highlight noteworthy findings.
We analyze how LMs and tasks are considered jointly in Figure 5. Despite the broad coverage, single studies, including fundamental ones, maintain a particular focus and consider only a fraction of LMs and tasks. For example, while most tasks (72%) were assessed on BERT, RoBERTa’s coverage has already declined to 42%. Conversely, part-of-speech tagging (POS), the most probed task, was only evaluated on 23% of the LMs, excluding prominent examples like BART Lewis et al. (2020). In particular, more recently released larger and powerful LMs, like Pythia (Biderman et al., 2023), UL2 (Tay et al., 2023), or LLAMA-2 (Touvron et al., 2023), as well as instruction-tuned LMs like FLAN-T5 (Chung et al., 2022) or LLAMA-2-Chat (Touvron et al., 2023) are missing almost entirely, with only a few recent exceptions (Hu and Levy, 2023; Waldis et al., 2024a). Again, these insights underscore the need to consolidate existing resources for more comprehensive coverage.
Figure 6 further highlights this point by sorting LMs and tasks according to their frequency of mention in relevant works and plotting their cumulative coverage. For example, considering all studies (red line), the top-10 most mentioned LMs account for 80% of all LMs mentions (black dot), while the remaining 151 unique LMs account for only 40%. A comparison of the paper’s focus reveals that methodological studies rely only on a limited set of 24 LMs and 36 tasks. In contrast, task-focused and analytical work cover a similar number of LMs (91 and 99, respectively). However, due to their distinct focus, task-focused studies cover a significantly larger number of tasks (202) than analytical ones (115).
@@
Type Phenomena Example Label
Morphology Subject-Verb Agreement And then, the cucumber was hurled into the air. Correct
And then, the cucumber were hurled into the air. Wrong
Syntax Part-of-Speech And then, the cucumber was hurled into the air. NN (Noun Singular)
Semantic Semantic Roles And then, the cucumber was hurled into the air. Direction
Reasoning Negation And then, the cucumber was hurled into the air. No Negation
Discourse Node Type in Rhetorical Tree And then, the cucumber was hurled into the air. Satellite
Table 1: Example instance of Holmes datasets for every type of linguistic phenomena. The relevant part of the example for the specific label is underlined.